Round to 4th decimal place, if needed A data set is as the following 8 1012 12 1214 14 141414 16 16 18 18 20 20 20 22 30 35 What is the mean? 2.What is the median? What is the variance, if this is a sample? What is the variance, if this is a population? Does it skew? (Right, Left, or Not) 5. A sample set is as the following with a missing value x110 12 12 12 1414 14 14 14 16 16 18 18 20 20 20 22 30 35 1. It is known that 18 What is the missing data value of the data set? What is the median? What is the variance? What is the variance, if this is a population? Does it skew? (Right, Left, or Not) 10. A population is listed as the followin 10 15 19 11. 12. 13. 14. 15. 16. Find 1 :? Find Σ(x-A)-? Find Σ(x-μ)-? Find Σ x =? Find Σ x2 =? Find σ-?A population is listed as the following 43 15747 Find the 2x2 Find the (Σ x)2 3 48 40 18 45 49 17. 18. 19. Find the 2x2 - 20. What is the standard deviation of the population?
Answer #1

1. Mean = Sum of all the data / Count of the data

= 8+10+12+12+12+14+14+14+14+14+16+16+18+18+20+20+20+22+30+35 / 20

= 339/20 = 16.95

2. Median = 11th data + 10th data / 2

= 14+16 / 2 = 30/2 = 15

3. Sample Variance = The average of the squared differences from the Mean.

= (16.95-8)2 + (16.95-10)2 and so on for all 20 values / 20

= 917.19/20 = 45.8595

4. Population Variance = Same as sample variance but because of unbiased estimator it will be divided by n-1 = 20-1 = 19

= (16.95-8)2 + (16.95-10)2 and so on for all 20 values / 19

= 917.19/19 = 48.2731

5. Yes, it is skewed rightwards, as we can see the last two data points are way higher than mean and also if you plot the graph, it will be skewed.

HOMEWORKLIB POLICY is to answer 1 question at a time so I can help you with others if you post them individually. Also, Every question can be solved now because all are same as first question. Thank You